OpenStudy (anonymous):

Find the center (h,k) and radius (r) of the circle: 2x^2+20x_2y^2=0

OpenStudy (anonymous):

** 2x^2 +20x+2y^2=0

OpenStudy (anonymous):

do you know what the general equation for a circle is?

OpenStudy (anonymous):

no not really, I missed this chapter of class

OpenStudy (anonymous):

\[\it is: (x-h)^{2} + (y-k)^{2} = r^{2}\] the point (h, k) is the circle's center. and r is the radius

OpenStudy (anonymous):

to put it in this form you'll need to complete the square for x. completing the square: \[x^{2} + bx = (x+ (b/2))^{2} - (b/2)^{2}\]

OpenStudy (anonymous):

so \[2x^{2} + 20x = 2(x^{2} + 10x) = 2[ (x + 5)^{2} - 25 ]\]

OpenStudy (anonymous):

we now have: \[2(x + 5)^{2} + 2y^{2} = 50\]

OpenStudy (anonymous):

\[= (x + 5)^{2} + y^{2} = 25\]

OpenStudy (anonymous):

where did the 2 go

OpenStudy (anonymous):

i divided both sides by two. it is equivalent from the general equation, can you tell me what the center and radius are? let me know when you want me to tell you

OpenStudy (anonymous):

(h,k) = (5,1) r = 25?

OpenStudy (anonymous):

(h, k) = (-5,0) r = sqrt25 = 5

OpenStudy (anonymous):

why is it the square root and not 25^2?

OpenStudy (anonymous):

the general equation has r^2 on the right side. meaning that r is really the sqrt of r^2

OpenStudy (anonymous):

oh ok! So, to graph this I would put a point at (-5, 0) and go out 5 spaces ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

proof : http://www.wolframalpha.com/input/?i=2%28x%2B5%29^2+%2B+2y^2+%3D+50 glad i could help :)

OpenStudy (anonymous):

so, if it touches (0,0) on the graph would that be an intercept or would (-10,0) be the only one?

OpenStudy (anonymous):

the point (0,0) would be considered both an x and y intercept

OpenStudy (anonymous):

Ok, so when I graph it and it asks for the intercepts I can say (-10, 0) and (0,0)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay :) thank you so much

OpenStudy (anonymous):

glad i could help :)